# Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model in the GENIE neutrino event generator

### v1

@@ III.1 Validation and comparison of models
For validation purposes, the model predictions obtained via GENIE have been compared with the original calculations performed by the authors of the model Martini et al. (2009), including the relativistic corrections for the quasielastic, published in Ref. Martini et al. (2011) for neutrinos and in Ref. Martini and Ericson (2013) for antineutrinos. The corresponding charged-current cross sections on 12C as a function of (anti)neutrino energy are presented for and in Figs. 2 and 2, respectively (and in Figs. 10 and 10 of Appendix References for 16O and 40Ca). These plots demonstrate perfect agreement between the original calculations (shaded bands) and the GENIE predictions for the different contributions to the cross section: “1p1h” (dashed), “npnh” (dotted) and their sum (solid). The predictions of other 1p1h and 2p2h models available in GENIE (“Nieves et al.” and “SuSAv2”) are also shown for reference. The corresponding GENIE tunes used are G18_10a_00_000 for the Nieves et al. model and G21_11a_00_000 for SuSAv2. As already illustrated in other articles (see for example Fig. 9 of Ref. Abe et al. (2023a)) the npnh (as well as the 1p1h) contributions estimated by these three models are different, and their relative importance with respect to 1p1h varies between neutrinos and antineutrinos, as discussed in Refs. Martini et al. (2010); Katori and Martini (2018). It is also important to note that a one-to-one correspondence between different exclusive channel contributions can be misleading. For example the nucleon-nucleon short-range correlations are part of npnh channel in the Martini et al. model, whereas they are included in the 1p1h component for SuSAv2. As can be seen in Figs. 2 and 2, the sum of these two channels is similar for Martini et al. and SuSAv2, especially in the antineutrino case.
@@
For the sake of comparison with other models, we show in Figs. 4 and 4 the T2K flux-integrated double differential cross section as a function of and for the 1p1h and npnh channels, corresponding to Fig. 4 of Dolan et al. (2022) and Fig. 2 of Dolan et al. (2020) respectively. In Fig. 4, one can easily recognize the quasielastic line
@@
where denotes the nucleon mass. The spread around this curve is due to Fermi motion111The fact that the quasielastic region is not strictly limited by the two quasielastic bands , typical of a global relativistic Fermi gas-based model with Fermi momentum , is due to the local density approximation and to the presence of Bessel functions in the nuclear response definitions (see Appendix 2 of Ref. Martini et al. (2009)), which result in very small fluctuations beyond the quasielastic region.. From Fig. 4 one can observe that the multinucleon response region covers the whole available phase space from the input tensors. Furthermore, two distinct branches can be identified: one below the quasielastic region – absent in the SuSAv2 case, see Fig. 2 of Ref. Dolan et al. (2020) – arising from nucleon-nucleon correlations, and a second one above the quasielastic region, extending from the dip to the region, due to -MEC contributions222These two branches are more distinguishable looking directly to the phase space of the nuclear response function, rather than to the flux-integrated cross section. See for example Fig. 7 of Ref. Katori and Martini (2018).. Overall, these two branches are less clearly separated than in the Nieves et al. model (see again Fig. 2 of Ref. Dolan et al. (2020)); in the Martini et al. case, the cross section is concentrated between the two contributions, which add coherently together with their significant interference. To summarize, the phase-space behavior of the multinucleon cross section in the Martini et al. model differs from those of Nieves et al. and SuSAv2, though it shares features with both: the presence of two branches, as in Nieves et al., and the accumulation of cross section strength in the dip region, as in SuSAv2.
@@
As a validation of the implementation, and to facilitate comparisons with other models, in Fig. 5 we present the double differential cross sections 333Here, denotes the solid angle element. Assuming azimuthal symmetry, the double differential cross section is related by . for both neutrino and antineutrino interactions on oxygen, at fixed incident energy and lepton scattering angle444GENIE does not allow the user to explicitly fix the scattering angle. Events with were selected to approximate a scattering angle of ., as a function of the energy transfer . This represents the most stringent validation test, as the cross sections are not integrated over any variable, in contrast to the total or flux-integrated differential cross sections discussed earlier. A perfect agreement is observed between the theoretical calculations (solid lines) and the GENIE implementation (histograms) for both 1p1h and npnh channels, and for neutrino as well as antineutrino interactions555In the 1p1h case, the small fluctuations at low discussed in footnote 1 can also be seen..
@@
Fig. 5 can be directly compared with Fig. 9 of Ref. Sobczyk and Nieves (2025), the recent reexamination by Sobczyk and Nieves of the multinucleon cross-section model originally presented in Ref. Nieves et al. (2011). In Ref. Sobczyk and Nieves (2025), the revised npnh single differential cross section at fixed neutrino energy, as well as the MiniBooNE flux-integrated double differential cross section on carbon, are compared with the corresponding results of Martini et al. Martini et al. (2009, 2011). The updated Valencia predictions are shown to be closer to the Lyon model, thus ending a long-standing discrepancy between the two.
@@
We now extend the comparison to the double differential cross section at fixed neutrino energy. Before focusing on the npnh channel, we note that the 1p1h distribution in Fig. 5 exhibits a shape similar to that in Fig. 9 of Ref. Sobczyk and Nieves (2025), as expected, since both are based on RPA calculations on top of a local relativistic Fermi gas. However, the distribution in Fig. 5 is slightly more peaked and shifted to lower by approximately . This is probably due to the absence of removal energy in the calculations of Martini et al. unlike the case of Nieves et al. These small differences are often washed out in flux-integrated double differential distributions (see, e.g., Fig. 10 of Ref. Sobczyk and Nieves (2025)).
@@ III.2 Limitations of the model implementation
Whilst the implementation described here represents a complete inclusion of the content of the Martini et al. model within the GENIE event generator for a set of defined nuclear targets, there are limitations that are important to note. Some of these are related to approximations inherent in producing a fully exclusive simulations from an inclusive input model, which are common to many model implementations in neutrino interaction generators. Others are due to missing predictive power within the Martini et al. model itself. Some of these limitations can be mitigated with relatively modest extensions of the implementation.
Firstly, it is important to note that the hadron tensors for both 1p1h and npnh interactions currently only go up to in energy transfer and in momentum transfer. This is due to a choice made in the original implementation of the model Martini et al. (2009). In the current GENIE implementation, outside of this range, the differential cross section vanishes (like for the other models following the hadron tensor implementation method). Further tractable but non-trivial model development work would be required to overcome this limitation. In principle, the implementation can be artificially expanded by interpolating (“blending”) the cross sections predicted by the Martini et al. model with those from another inclusive model, such as SuSAv2 (as already done within the CRPA model implementation scheme in GENIE Dolan et al. (2022)). For the purposes of this implementation, as well as the comparisons with flux-averaged measurements we consider in this work, the current implementation is sufficient.
@@
We additionally note that, although the implementation features a 3p3h contribution to the npnh channel, this has been added as an additional “effective” 2p2h channel. This means that only two nucleons are produced at the vertex for a 3p3h interaction, relying solely on the existing GENIE hadron kinematics generator. This could be mitigated in a future extension to this implementation to produce a third final-state nucleon (via a new interaction channel within GENIE)666Separated 2p2h and 3p3h hadron tensors are already available.. Again, predictions in this article are either entirely unaffected or not very sensitive to alterations to the treatment of the hadronic system.
@@
The current implementation relies on hadron tensors computed for three isoscalar nuclei – 12C, 16O and 40Ca. For nuclear targets other than those for which the tensors have been calculated, the total 1p1h cross section is scaled according to the difference in the number of target nucleons (i.e. depending on whether the incoming probe is a neutrino or an antineutrino). For example, for an 40Ar target, the total 1p1h cross section on 40Ca is scaled by a factor of 22/20 (18/20) for an incoming neutrino (antineutrino) probe. For npnh interactions, the scaling of the total cross section to different targets is based on atomic number, also following the scheme in Ref. Dolan et al. (2020). Note that a small additional scaling is applied, following the same scheme as for the SuSAv2 model Dolan et al. (2020) based on superscaling argument. This is the (square of) ratio of the Fermi momenta of the reference and target nuclei for 1p1h (npnh) interactions. Whilst not perfect, this overall scaling is expected to be a reasonable approximation away from the low energy transfer region (100 MeV).
@@
More generally, the Martini et al. model internally does not differentiate between protons and neutrons when calculating cross sections, meaning that the impact of non-isoscalarity of target nuclei is not accounted for. Mitigating this limitation can be a challenging task. A first step could be assigning different Fermi momenta () to neutrons and protons within the model. However, this naive scaling does not take into account the modification of RPA suppression for asymmetric nuclei.
@@
Finally, the Martini et al. model, and therefore the hadron tensors implemented in GENIE, do not take into account the impact of the nuclear removal energy. The consequence of this is that the lepton kinematics, in both GENIE and the Martini et al. model, are calculated assuming an interaction with unbound (but still moving) nucleons. On the other hand, GENIE does consider the removal energy when calculating hadron kinematics via the aforementioned factorization scheme. Whilst in this paper we do not attempt to modify the original inclusive predictions of the Martini et al. model, a possible mitigation for this limitation could be an extension of this implementation to add an effective modification to the energy transfer read from the hadron tensors to simulate the impact of removal energy. This strategy can be trivially adopted using routines already available within the GENIE generator.
@@
Overall, the model implementation is a faithful reproduction of the Martini et al. model for outgoing lepton kinematics for carbon, oxygen or calcium nuclei. Scaling to different nuclei of similar atomic number is broadly reliable. The hadron kinematics predicted by the GENIE from this model implementation should be treated with cautious skepticism (as is the case also for other 1p1h and 2p2h GENIE model implementations).
@@ IV Comparison with measurements
As discussed in Sec. II, the theoretical calculations by Martini et al. have already been compared in the past with several cross-section measurements. However, previous comparisons were performed in the absence of a unified generator implementation, so the treatment of effects such as final-state interactions (FSI) was not applied in a consistent way to all interaction channels. Here we present a few additional comparisons to illustrate the potential of the present GENIE implementation to make cross-section predictions for different neutrino beams, various nuclear targets and diverse final-state topologies, using a unified generator framework as enabled by this implementation.
@@
To this end, we show the T2K flux-integrated charged-current double differential cross sections without pions in the final state, i.e. the “CC0” cross section, on carbon (Fig. 7(a)) and oxygen (Fig. 7(b)). The experimental T2K results are the regularized cross sections obtained via simultaneous analysis of the carbon and oxygen target samples, see Ref. Abe et al. (2020c) (and Fig. 9 in this publication) for more details. To process the simulated events, which include pion absorption contribution as implemented in GENIE, we use the NUISANCE framework Stowell et al. (2017) which facilitates direct measurement-model comparisons.
@@
The Martini et al. model yields a reasonable agreement with the T2K measurement. The total for both carbon and oxygen distributions is for 58 degrees of freedom777A similar comparison with the so-called unregularized cross sections Abe et al. (2020c) yields for 58 degrees of freedom., which corresponds to a p-value of 0.05, indicating satisfactory agreement. This result can be compared with the values presented in Table V of Ref. Abe et al. (2020c) which presents the for nine different versions/tunes of the neutrino event generators888Note that some of these models have been updated since the measurement publication.. The obtained with the Martini et al. model is lower than that of five of these models and higher than that of four. It is useful to note that Table V of Ref. Abe et al. (2020c) emphasizes that the majority of the discrepancy between the models and the measurement is driven by the most forward bin, which is dominated by low energy transfer interactions. This is the region where suppression of the cross section from long-range correlations via RPA are most pronounced.
@@
The Martini et al. model generally performs better than both the NEUT and NuWro Spectral Function (SF) models. This is expected, as these model implementations do not account for suppression of the cross section via RPA effects. The Martini et al. model also performs better than the GENIE SuSAv2 implementation and the RMF 1p1h+SuSAv2 prediction. In the first case, the SuSAv2 1p1h model implementation in GENIE does not include an explicit RPA suppression (but does include an effective suppression of the cross section at low energy transfer due to FSI, for more details see Dolan et al. (2020)), which results in an overprediction of the measurement in the most forward bins. In the case of the RMF model, the effect of FSI is included in the distortion of the final wavefunction from first principles, which causes a larger reduction in the cross section and thus a better agreement with the measurement. The Martini et al. model performs better than both of these models, indicating the probable need for an even larger suppression due to RPA effects. Finally, our implementation yields a lower than the GiBUU model tested in Ref. Abe et al. (2020c). This may be due to the fact that the LFG model used in GiBUU Buss et al. (2012) does not apply any RPA effects.
@@
Conversely, the Martini et al. model yields a higher than the other GENIE, and NEUT Hayato and Pickering (2021) LFG+RPA-based models. In both cases, these implementations are based on the model of Nieves et al. Nieves et al. (2011) and, as discussed in Sec. III.1, have similar predictions for the inclusive cross section on carbon to those of the Martini et al. model (see Fig. 2 and Fig. 9 of Ref. Sobczyk and Nieves (2025)). A significant portion of the difference in the between the LFG-based models and the Martini et al. is therefore likely due to the treatment of 2p2h interactions. Whereas the updated treatment of multi-nucleon interactions discussed in Ref. Sobczyk and Nieves (2025) reduces the disagreement between the Martini et al. and the Nieves et al. models, the treatment used in the comparisons of Ref. Abe et al. (2020c) predicted a 2p2h cross section for the Nieves et al. model which is significantly lower than that of the Martini et al. (as shown in Fig. 2). It is additionally worth noting that the quoted values are for the combined carbon and oxygen measurement. A closer inspection of the individual carbon and oxygen inclusive predictions in the most forward bin reveals that the Nieves et al. and Martini et al. predictions are extremely close for carbon, but they differ slightly for oxygen. This implies that some missing ingredients in the detailed differences between the cross sections on carbon and oxygen affecting the forward region (for example concerning the nuclear removal energy) may be responsible for the higher .
@@
Another example of model-measurement comparison is shown in Fig. 8 which presents the MicroBooNE flux integrated charged-current single differential cross sections without pions and with one proton in the final state – the so called “CC1p0” cross section – on argon. Compared to the previous T2K case, this measurement involves a different neutrino flux, a different nuclear target, and a different final-state topology. The measurement is that published by MicroBooNE in Ref. Abratenko et al. (2023), where the signal includes all -Ar scattering events with a final-state muon of momentum and exactly one proton with . Regarding the modeling, we recall that the implemented microscopic calculations are those of 40Ca, rescaled to 40Ar by following the same procedure discussed and justified in Ref. Martini et al. (2022), to which we refer the reader for details. Fig. 8 corresponds to Fig. 11 of Ref. Abratenko et al. (2023), which includes comparisons with other model predictions and the corresponding . In this case, the obtained with the Martini et al. model is for 18 degrees of freedom, which corresponds to a p-value of 0.994. Among the models considered, the Martini et al. model provides the lowest value for this particular measurement. This is likely due to the fact that the 2p2h cross section predicted by our implementation is significantly higher than that predicted by all of the other models considered in Fig. 11 of Ref. Abratenko et al. (2023), which are based on the previous version of the Nieves et al. model. Other studies (e.g. Ref. Filali et al. (2025)) have indeed shown that more exclusive MicroBooNE measurements indicate a preference for an enhanced 2p2h contribution on argon.

### v2

@@ III.1 Validation and comparison of models
For validation purposes, the model predictions obtained via GENIE have been compared with the original calculations performed by the authors of the model Martini et al. (2009), including the relativistic corrections for the quasielastic, published in Ref. Martini et al. (2011) for neutrinos and in Ref. Martini and Ericson (2013) for antineutrinos. The corresponding charged-current cross sections on 12C as a function of (anti)neutrino energy are presented for and in Figs. 2 and 2, respectively (and in Figs. 10 and 10 of Appendix A for 16O and 40Ca). These plots demonstrate perfect agreement between the original calculations (shaded bands) and the GENIE predictions for the different contributions to the cross section: “1p1h” (dashed), “npnh” (dotted) and their sum (solid). The predictions of other 1p1h and 2p2h models available in GENIE (Nieves et al. and SuSAv2) are also shown for reference. The corresponding GENIE tunes used are G18_10a_00_000 for the Nieves et al. model and G21_11a_00_000 for SuSAv2. As already illustrated in other articles (see, for example, Fig. 9 of Ref. Abe et al. (2023a)), the npnh (as well as the 1p1h) contributions estimated by these three models are different, and their relative importance with respect to 1p1h varies between neutrinos and antineutrinos, as discussed in Refs. Martini et al. (2010); Katori and Martini (2018). It is also important to note that a one-to-one correspondence between different exclusive channel contributions can be misleading. For example the nucleon-nucleon short-range correlations are part of npnh channel in the Martini et al. model, whereas they are included in the 1p1h component for SuSAv2. As can be seen in Figs. 2 and 2, the sum of these two channels is similar for Martini et al. and SuSAv2, especially in the antineutrino case.
@@
For the sake of comparison with other models, we show in Figs. 4 and 4 the T2K flux-integrated double differential cross section as a function of and for the 1p1h and npnh channels, corresponding to Fig. 4 of Dolan et al. (2022) and Fig. 2 of Dolan et al. (2020), respectively. In Fig. 4, one can easily recognize the quasielastic line
@@
where denotes the nucleon mass. The spread around this curve is due to Fermi motion333The fact that the quasielastic region is not strictly limited by the two quasielastic bands , typical of a global relativistic Fermi gas-based model with Fermi momentum , is due to the local density approximation and to the presence of Bessel functions in the nuclear response definitions (see Appendix 2 of Ref. Martini et al. (2009)), which result in very small fluctuations beyond the quasielastic region.. From Fig. 4 one can observe that the multinucleon response region covers the whole available phase space from the input tensors. Furthermore, two distinct branches can be identified: one below the quasielastic region–absent in the SuSAv2 case, see Fig. 2 of Ref. Dolan et al. (2020)–arising from nucleon-nucleon correlations, and a second one above the quasielastic region, extending from the dip to the region, due to -MEC contributions444These two branches are more distinguishable looking directly to the phase space of the nuclear response function, rather than to the flux-integrated cross section. See for example Fig. 7 of Ref. Katori and Martini (2018).. Overall, these two branches are less clearly separated than in the Nieves et al. model (see again Fig. 2 of Ref. Dolan et al. (2020)); in the Martini et al. case, the cross section is concentrated between the two contributions, which add coherently together with their significant interference. To summarize, the phase-space behavior of the multinucleon cross section in the Martini et al. model differs from those of Nieves et al. and SuSAv2, though it shares features with both: the presence of two branches, as in Nieves et al., and the accumulation of cross-section strength in the dip region, as in SuSAv2.
@@
As a validation of the implementation, and to facilitate comparisons with other models, in Fig. 5 we present the double differential cross sections555Here, denotes the solid angle element. Assuming azimuthal symmetry, the double differential cross section is related by . for both neutrino and antineutrino interactions on oxygen, at fixed incident energy and lepton scattering angle,666GENIE does not allow the user to explicitly fix the scattering angle. Events with were selected to approximate a scattering angle of . as a function of the energy transfer . This represents the most stringent validation test, as the cross sections are not integrated over any variable, in contrast to the total or flux-integrated differential cross sections discussed earlier. A perfect agreement is observed between the theoretical calculations (solid lines) and the GENIE implementation (histograms) for both 1p1h and npnh channels, and for neutrino as well as antineutrino interactions.777In the 1p1h case, the small fluctuations at low discussed in footnote 3 can also be seen.
@@
Figure 5 can be directly compared with Fig. 9 of Ref. Sobczyk and Nieves (2025), the recent reexamination by Sobczyk and Nieves of the multinucleon cross-section model originally presented in Ref. Nieves et al. (2011). In Ref. Sobczyk and Nieves (2025), the revised npnh single differential cross section at fixed neutrino energy, as well as the MiniBooNE flux-integrated double differential cross section on carbon, are compared with the corresponding results of Martini et al. Martini et al. (2009, 2011). The updated Valencia predictions are shown to be closer to the Lyon model, thus ending a long-standing discrepancy between the two.
@@
We now extend the comparison to the double differential cross section at fixed neutrino energy. Before focusing on the npnh channel, we note that the 1p1h distribution in Fig. 5 exhibits a shape similar to that in Fig. 9 of Ref. Sobczyk and Nieves (2025), as expected, since both are based on RPA calculations on top of a local relativistic Fermi gas. However, the distribution in Fig. 5 is slightly more peaked and shifted to lower by approximately . This is probably due to the absence of removal energy in the calculations of Martini et al., unlike the case of Nieves et al. These small differences are often washed out in flux-integrated double differential distributions (see, e.g., Fig. 10 of Ref. Sobczyk and Nieves (2025)).
@@ III.2 Limitations of the model implementation
While the implementation described here represents an inclusion of the 1p1h and npnh channels of the Martini et al. model within the GENIE event generator for a set of defined nuclear targets for charged-current neutrino and antineutrino scattering, there are limitations that are important to note. Some of these are related to approximations inherent in producing a fully exclusive simulations from an inclusive input model, which are common to many model implementations in neutrino interaction generators. Others are due to missing predictive power within the Martini et al. model itself. Some of these limitations can be mitigated with relatively modest extensions of the implementation.
First, it is important to note that the hadron tensors for both 1p1h and npnh interactions currently only go up to in energy transfer and in momentum transfer. This is due to a choice made in the original implementation of the model Martini et al. (2009). In the current GENIE implementation, outside of this range, the differential cross section vanishes (like for the other models following the hadron tensor implementation method). Further tractable but nontrivial model development work would be required to overcome this limitation. In principle, the implementation can be artificially expanded by interpolating (“blending”) the cross sections predicted by the Martini et al. model with those from another inclusive model, such as SuSAv2 (as already done within the CRPA model implementation scheme in GENIE Dolan et al. (2022)). For the purposes of this implementation, as well as the comparisons with flux-averaged measurements we consider in this work, the current implementation is sufficient. We remind the reader also that, even if the tensors start from , giant resonance peaks that may appear at low transferred momenta for are not included in the formalism of Martini et al. at difference of the CRPA Pandey et al. (2015); Dolan et al. (2022) approach, particularly suited for this low-energy low-momentum region. However, in this region, the RPA Martini et al. calculations (which do not include giant resonance peaks) may represent a reasonable average of the CRPA results Pandey et al. (2015), as shown in Fig. 2 of Ref. Martini et al. (2016) and discussed in that article.
@@
We additionally note that, although the implementation features a 3p3h contribution to the npnh channel, this has been added as an additional “effective” 2p2h channel. This means that only two nucleons are produced at the vertex for a 3p3h interaction, relying solely on the existing GENIE hadron kinematics generator. This could be mitigated in a future extension to this implementation to produce a third final-state nucleon (via a new interaction channel within GENIE)888We note that separate 2p2h and 3p3h hadron tensors are available if the user prefers to only simulate one of the two npnh contributions, or a specific sub-contribution within the 2p2h channel as described in the final paragraph of Sec. II. This option can be configured by editing the config/MartiniMECHadronTensorModel.xml configuration file.. This was not implemented at the time of writing this article as the current implementation uses the same routine to process the combined npnh (i.e. 2p2h+3p3h) tensors. Separating the two types of interactions would have required creating a new routine for the 3p3h channel. Since the current model predictions are inclusive, we leave this extension for future work. Again, predictions in this article are either entirely unaffected or not very sensitive to alterations to the treatment of the hadronic system.
@@
The current implementation relies on hadron tensors computed for three isoscalar nuclei–12C, 16O and 40Ca. For nuclear targets other than those for which the tensors have been calculated, the total 1p1h cross section is scaled according to the difference in the number of target nucleons (i.e. depending on whether the incoming probe is a neutrino or an antineutrino). For example, for an 40Ar target, the total 1p1h cross section on 40Ca is scaled by a factor of 22/20 (18/20) for an incoming neutrino (antineutrino) probe. For npnh interactions, the total cross section is scaled with the mass number, also following the scheme in Ref. Dolan et al. (2020), implying no scaling between the 40Ca and 40Ar npnh cross sections. Note that a small additional scaling is applied, following the same scheme as for the SuSAv2 model Dolan et al. (2020) based on superscaling arguments. This is the (square of the) ratio of the Fermi momenta of the reference and target nuclei for 1p1h (npnh) interactions.999The Fermi momentum used for this calculation are taken from Ref. Maieron et al. (2002). It is assumed to be the same for protons and neutrons. For nuclei not in the table the Fermi momentum of the nucleus with the closest proton number is used. While not perfect, this overall scaling is expected to be a reasonable approximation away from the low-energy transfer region (100 MeV). Scaling to heavier nuclear targets occasionally employed in neutrino scattering measurements, such as iron or lead, could be improved with a future addition of a hadron tensor table for a heavier target.
@@
More generally, the Martini et al. model internally does not differentiate between protons and neutrons when calculating cross sections, meaning that the impact of nonisoscalarity of target nuclei is not accounted for. Mitigating this limitation can be a challenging task. A first step could be assigning different Fermi momenta () to neutrons and protons within the model. However, this naive scaling does not take into account the RPA modification for asymmetric nuclei Alberico et al. (1989), a challenge that, to our knowledge, is shared with the other RPA-based models at present, theoretical calculations of asymmetric relativistic Fermi gas for quasielastic lepton-nucleus scattering being only relatively recent Barbaro et al. (2018). It is also important to stress that the first microscopic calculation of two-particle-two-hole meson-exchange currents in 40Ar discussing also asymmetric scaling properties for neutrino and electron scattering appeared only very recently Martinez-Consentino et al. (2025).
@@
Finally, the Martini et al. model, and therefore the hadron tensors implemented in GENIE, do not take into account the impact of the nuclear removal energy. The consequence of this is that the lepton kinematics, in both GENIE and the Martini et al. model, are calculated assuming an interaction with unbound (but still moving) nucleons. On the other hand, GENIE does consider the removal energy when calculating hadron kinematics via the aforementioned factorization scheme. While in this paper we do not attempt to modify the original inclusive predictions of the Martini et al. model, a possible mitigation for this limitation could be an extension of this implementation to add an effective modification to the energy transfer read from the hadron tensors to simulate the impact of removal energy. This strategy can be trivially adopted using routines already available within the GENIE generator.
@@
Overall, the model implementation is a faithful reproduction of the Martini et al. model for outgoing lepton kinematics for carbon, oxygen or calcium nuclei. Scaling to different nuclei of similar atomic number is broadly reliable. The hadron kinematics predicted by GENIE from this model implementation should be treated with cautious skepticism (as is the case also for other 1p1h and 2p2h GENIE model implementations).
@@ IV Comparison with measurements
As discussed in Sec. II, the theoretical calculations by Martini et al. have already been compared in the past with several cross-section measurements. However, previous comparisons were performed in the absence of a unified generator implementation, so the treatment of effects such as final-state interactions (FSIs) was not applied in a consistent way to all interaction channels. Here we present a few additional comparisons to illustrate the potential of the present GENIE implementation to make cross-section predictions for different neutrino beams, various nuclear targets and diverse final-state topologies, using a unified generator framework as enabled by this implementation.
Throughout this section, we compare to two experimental measurements from the T2K and MicroBooNE Collaborations, both of which report the cross section for the so-called CC topology. This topology contains contributions from the 1p1h and npnh channels, which are simulated with the Martini et al. model implementation described in this article. There is also a subleading contribution from interactions where a pion was produced (usually via a resonant interaction) and then absorbed via FSI. To model this contribution, we use the standard GENIE prescription (Ref. Alvarez-Ruso et al. (2021)) which employs the Berger-Sehgal model (Ref. Berger and Sehgal (2007)) with updated form factors from Graczyk and Sobczyk (Ref. Graczyk and Sobczyk (2008)) and includes a variety of resonances up to invariant hadronic mass GeV, but neglecting interferences between them. Final-state interactions are simulated via the GENIE intranuclear cascade using the hA2018Intranuke routine (for details, see Ref. Dytman et al. (2021)). FSIs will impact the predictions in several ways. Although the FSI model in GENIE does not modify the total inclusive cross section, it will affect the hadrons that are used in defining the signal topologies. For both T2K and MicroBooNE, FSI impacts the samples mainly via pion absorption, which contributes to the measured signal sample. In addition to this, the MicroBooNE samples require at least one proton with momentum above , and the final-state proton kinematics that determine the sample composition will also be impacted by the choice of the FSI model (its effect is notably to decelerate protons and thus reduce the total cross section). It is also worth noting that the MicroBooNE samples also allow final-state charged pions with momenta below , which can also be affected by the choice of resonant pion production and FSI models. Finally, we note that FSI can also impact both T2K and MicroBooNE samples via pion production from nucleon interactions with the residual nucleus, but the contribution of these processes in this energy range is negligible for the purposes of this discussion.
@@
We begin by showing the T2K flux-integrated charged-current double differential cross sections without pions in the final state, i.e. the “CC0” cross section, on carbon (Fig. 7(a)) and oxygen (Fig. 7(b)). The experimental T2K results are the regularized cross sections obtained via simultaneous analysis of the carbon and oxygen target samples, see Ref. Abe et al. (2020c) (and Fig. 9 in this publication) for more details. To process the simulated events, which include pion absorption contribution as implemented in GENIE, we use the NUISANCE framework Stowell et al. (2017) which facilitates direct measurement-model comparisons.
@@
The Martini et al. model yields a reasonable agreement with the T2K measurement. The total for both carbon and oxygen distributions is for 58 degrees of freedom,101010A similar comparison with the so-called unregularized cross sections Abe et al. (2020c) yields for 58 degrees of freedom. which corresponds to a p value of 0.05, indicating satisfactory agreement. This result can be compared with the values presented in Table V of Ref. Abe et al. (2020c) which presents the for nine different versions/tunes of the neutrino event generators.111111Note that some of these models have been updated since the measurement publication. The obtained with the Martini et al. model is lower than that of five of these models and higher than that of four. It is useful to note that Table V of Ref. Abe et al. (2020c) emphasizes that the majority of the discrepancy between the models and the measurement is driven by the most forward bin, which is dominated by low-energy transfer interactions. This is the region where suppression of the cross section from long-range correlations via RPA is most pronounced.
@@
The Martini et al. model generally performs better than both the NEUT and NuWro spectral function models. This is expected, as these model implementations do not account for suppression of the cross section via RPA effects. The Martini et al. model also performs better than the GENIE SuSAv2 implementation and the RMF 1p1h+SuSAv2 prediction. In the first case, the SuSAv2 1p1h model implementation in GENIE does not include an explicit RPA suppression (but does include an effective suppression of the cross section at low-energy transfer due to FSI; for more details, see Dolan et al. (2020)), which results in an overprediction of the measurement in the most forward cos bins. In the case of the RMF model, the effect of FSI is included in the distortion of the final wave function from first principles, which causes a larger reduction in the cross section and thus a better agreement with the measurement. The Martini et al. model performs better than both of these models, indicating the probable need for an even larger suppression due to RPA effects. Finally, our implementation yields a lower than the GiBUU model tested in Ref. Abe et al. (2020c). This may be due to the fact that the LFG model used in GiBUU Buss et al. (2012) does not apply any RPA effects.
@@
Conversely, the Martini et al. model yields a higher than the other GENIE and NEUT Hayato and Pickering (2021) LFG+RPA-based models. In both cases, these implementations are based on the model of Nieves et al. Nieves et al. (2011) and, as discussed in Sec. III.1, have similar predictions for the inclusive cross section on carbon to those of the Martini et al. model (see Figs. 2 and 9 of Ref. Sobczyk and Nieves (2025)). A significant portion of the difference in the between the LFG-based models and the Martini et al. model is therefore likely due to the treatment of 2p2h interactions. Whereas the updated treatment of multinucleon interactions discussed in Ref. Sobczyk and Nieves (2025) reduces the disagreement between the Martini et al. and the Nieves et al. models, the treatment used in the comparisons of Ref. Abe et al. (2020c) predicted a 2p2h cross section for the Nieves et al. model which is significantly lower than that of the Martini et al. (as shown in Fig. 2). It is additionally worth noting that the quoted values are for the combined carbon and oxygen measurement. A closer inspection of the individual carbon and oxygen inclusive predictions in the most forward bin reveals that the Nieves et al. and Martini et al. predictions are extremely close for carbon, but they differ slightly for oxygen. This implies that some missing ingredients in the detailed differences between the cross sections on carbon and oxygen affecting the forward region (for example concerning the nuclear removal energy) may be responsible for the higher .
@@
Another example of model-measurement comparison is shown in Fig. 8, which presents the MicroBooNE flux-integrated charged-current single differential cross sections without pions and with one proton in the final state–the so called CC1p0 cross section–on argon. Compared to the previous T2K case, this measurement involves a different neutrino flux, a different nuclear target, and a different final-state topology. The measurement is that published by MicroBooNE in Ref. Abratenko et al. (2023), where the signal includes all -Ar scattering events with a final-state muon of momentum and exactly one proton with . Regarding the modeling, we recall that the implemented microscopic calculations are those of 40Ca, rescaled to 40Ar by following the same procedure discussed and justified in Ref. Martini et al. (2022), to which we refer the reader for details. Figure 8 corresponds to Fig. 11 of Ref. Abratenko et al. (2023), which includes comparisons with other model predictions and the corresponding . In this case, the obtained with the Martini et al. model is for 18 degrees of freedom, which corresponds to a p value of 0.994. Among the models considered, the Martini et al. model provides the lowest value for this particular measurement. This is likely due to the fact that the 2p2h cross section predicted by our implementation is significantly higher than that predicted by all of the other models considered in Fig. 11 of Ref. Abratenko et al. (2023), which are all based on the previous version of the Nieves et al. model but GiBUU.121212We notice that in Fig. 11 of Ref. Abratenko et al. (2023) GiBUU provides the lowest () with respect to the other models considered in that figure. This reinforces our hypothesis that the reason is related to the 2p2h contribution, the GiBUU one being larger than the one of the other Monte Carlos, all based on the previous version of the Nieves et al. model, and slightly lower than the predictions of Martini et al. for argon, as one can appreciate by comparing Fig. 8 of Ref. Mosel and Gallmeister (2024) with Fig. 5 of Ref. Martini et al. (2022). Other studies (e.g., Ref. Filali et al. (2025)) have indeed shown that more exclusive MicroBooNE measurements indicate a preference for an enhanced 2p2h contribution on argon.